# Márton Benedek 1122    Péter Biró 2211    Gergely Csáji 22 Matthew Johnson 33

### v2

@@ 3 ILP Formulation
In the kidney exchange literature the following ILP is called the edge-formulation (see e.g. [1]). For each $a\in A$ in the graph $D$, let $e_{a}\in\{0,1\}$ be a (binary) edge-variable; this yields a vector $e$ with entries $e_{a}$.
$\begin{array}[]{rl}M^{*}:=\displaystyle\max_{e}\sum_{ij\in A}e_{ij}&\text{s.t.}\\[-5.0pt] \end{array}$ (edge-formulation)
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\ \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\ \end{array}$
The first set of constraints represents the (well-known) Kirchoff law. The second set of constraints ensures that every node is covered by at most one cycle. The objective function provides a maximum cycle packing (of size $M^{*}$).
@@
Once we have $M^{*}$ we solve the following ILP to find $d_{1}^{*}$:
$\begin{array}[]{rl}d_{1}^{*}:=\displaystyle\min_{e,d_{1}}d_{1}&\text{s.t.}\\ \end{array}$ ($\text{ILP}_{d_{1}}$)
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[18.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[18.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[18.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}&\forall p\in N\\[18.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}&\forall p\in N\end{array}$
The first three constraints guarantee that all solutions are, in fact, maximum cycle packings. These constraints will be part of the formulation throughout the entire ILP-series. The remaining two constraints, together with the objective function, guarantee that we minimize the largest country deviation. Note that for each country $p$, exactly one of $\sum_{j\in V_{p}}e_{ij}-x_{p}$ and $x_{p}-\sum_{j\in V_{p}}e_{ij}$ is positive and exactly one is negative, unless both are zero. However, as soon as we reach $d_{t}^{*}\leq 1/2$ we have found a strongly close maximum cycle packing. Hence, in the remainder, we assume $d_{t}^{*}>1/2$ for each $t$ such that the series continues with $t+1$.
($\text{ILP}_{d_{1}}$) has one additional continuous variable ($d_{1}$) and $12|N|+$ 1 additional constraints. For every country $p\in N$, we have that $|\sum_{j\in V_{p}}e_{ij}-x_{p}|\leq d_{1}^{*}$. However, there exists a smallest subset $N_{1}\subseteq N$ (which may not necessarily be unique) such that
$\begin{array}[]{ll}\displaystyle\left|\sum_{j\in V_{p}}e_{ij}-x_{p}\right|=d_{1}^{*}&\forall p\in N_{1}\\ &\\ \displaystyle\left|\sum_{j\in V_{p}}e_{ij}-x_{p}\right|\leq d_{2}^{*}<d_{1}^{*}&\forall p\in N\setminus N_{1}\end{array}$
@@
The number of vertices for a country covered by any cycle packing is an integer. Hence, the number of possible country deviations is at most $2|N|$ and depends only on $x$. The fractional part of a country deviation is either $\operatorname{frac}(x_{p})$ or $1-\operatorname{frac}(x_{p})$. Therefore, to find the minimal positive difference in between the deviations of any two countries $p$ and $q$, we have to compare the values $\operatorname{frac}(x_{p})$ and $1-\operatorname{frac}(x_{p})$, with $\operatorname{frac}(x_{q})$ or $1-\operatorname{frac}(x_{q})$ and take the minimum of those four possible differences. Let $\varepsilon$ be a small positive constant that is strictly smaller than the minimum possible positive difference between any two countries.
@@
$\begin{array}[]{rl}\displaystyle\min_{z^{1},e}\displaystyle\sum_{p\in N}z^{1}_{p}&\text{s.t.}\\ \end{array}$ ($\text{ILP}_{N_{1}}$)
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}-\varepsilon(1-z^{1}_{p})&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}-\varepsilon(1-z^{1}_{p})&\forall p\in N\end{array}$
@@
As discussed, for each country $p$, the left hand side of either the fourth or the fifth constraint is negative (i.e., would be satisfied even with $z^{1}_{p}=0$). For those countries whose deviation cannot be lower than $d_{1}^{*}$, however, the (positive) left hand side of either the fourth or the fifth constraint will require $z^{1}_{p}=1$. Thus, given an optimal solution $z^{1*}$ of ($\text{ILP}_{N_{1}}$), let $n_{1}^{*}:=\sum_{p\in N}z_{p}^{1*}$ be the minimal number of countries receiving the largest country deviations. It is guaranteed that the non-increasingly ordered country deviations at a strongly close maximum cycle packing starts with exactly $n_{1}^{*}$ many $d_{1}^{*}$ values, followed by some $d_{2}^{*}<d_{1}^{*}$. ($\text{ILP}_{N_{1}}$) has $|A|+|N|$ binary variables and $12|V|+2|N|+$ 1 constraints. Now, to find $d_{2}^{*}$, we solve the following ILP:
$\begin{array}[]{rl}\displaystyle\min_{d_{2},e,z^{1}}d_{2}&\text{s.t.}\\ \end{array}$ ($\text{ILP}_{d_{2}}$)
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{2}+z^{1}_{p}d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{2}+z^{1}_{p}d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{p\in N}z^{1}_{p}&=&n_{1}^{*}&\end{array}$
@@
($\text{ILP}_{d_{2}}$) has $|A|+|N|$ binary variables and one continuous variable ($d_{2}$) with $22|V|+4|N|+$ 2 constraints, and guarantees that we find the minimal second-largest country deviation $d_{2}^{*}$ while exactly $n_{1}^{*}$ countries deviation is kept at $d_{1}^{*}$. Finding $n_{2}^{*}$ follows a similar approach, where $L$ is a large constant satisfying $L\geq 2(d_{1}^{*}-d_{2}^{*})$:
$\begin{array}[]{rl}\displaystyle\min_{z^{1},z^{2},e}\displaystyle\sum_{p\in N}z^{2}_{p}&\text{s.t.}\\ \end{array}$ ($\text{ILP}_{N_{2}}$)
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{p\in N}z^{1}_{p}&=&n_{1}^{*}&\\[15.0pt] z_{p}^{1}+z_{p}^{2}&\leq&1&\forall p\in N\end{array}$
@@
$\begin{array}[]{rl}\displaystyle\min_{d_{t},e,(z^{i})_{i=1}^{t-1}}d_{t}&\text{s.t.}\\ \end{array}$ ($\text{ILP}_{d_{t}}$)
@@
In the following formulations $L$ is a large constant satisfying $L\geq d_{t}^{*}-\varepsilon$.
$\begin{array}[]{rl}\displaystyle\min_{(z)_{i=1}^{t},e}\displaystyle\sum_{p\in N}z^{t}_{p}&\text{s.t.}\\ \end{array}$ ($\text{ILP}_{N_{t}}$)
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{i=1}^{t}z^{i}_{p}&\leq&1&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}&\forall p\in N\\[15.0pt] x_{p}-\displaystyle\sum_{j\in V_{p}}e_{ij}&\leq&d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i})L&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i})L&\forall p\in N\\[20.0pt] \displaystyle\sum_{p\in N}z^{i}_{p}&=&n_{i}^{*}&\forall i=1,\dots,t-1\end{array}$
From the above, we conclude that the following theorem holds.
@@
For a partitioned permutation games $(N,v)$ defined on a graph $G=(V,A)$, it is possible to find an optimal solution that is strongly close to a given target allocation $x$ by solving a series of at most $2|N|$ ILPs, each having $O(|A|+|N|^{2})$ binary variables and $O(|V|)$ constraints.

### v3

@@ 3 ILP Formulation
In the kidney exchange literature the following ILP is called the edge-formulation (see e.g. [ABS07]). For each ${ij}\in A$ in the graph $G$, let $e_{{ij}}\in\{0,1\}$ be a (binary) edge-variable; this yields a vector $e$ with entries $e_{{ij}}$.
[…]
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $=$ $\displaystyle\sum_{j:ij\in A}e_{ij}$ $\quad\forall i\in V$ (1)
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $\leq$ $\quad 1$ $\quad\forall i\in V$ (2)
Constraint (1) represents the well-known Kirchoff law, whereas (2) ensures that every node is covered by at most one cycle. The objective function provides a maximum cycle packing, which has size $\mu^{*}$.
@@
Once we have $\mu^{*}$ we solve the following ILP to find $d_{1}^{*}$:
[…]
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $=$ $\displaystyle\sum_{j:ij\in A}e_{ij}$ $\quad\forall i\in V$ (1)
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $\leq$ $\quad 1$ $\quad\forall i\in V$ (2)
$\displaystyle\sum_{ij\in A}e_{ij}$ $=$ $\quad\mu^{*}$ (3)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad d_{1}$ $\quad\forall p\in N$ (4)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad d_{1}$ $\quad\forall p\in N$ (5)
Constraints (1)-(3) guarantee that all solutions are, in fact, maximum cycle packings. These constraints will be part of the formulations throughout the entire ILP-series. Constraints (4) and (5), together with the objective function, guarantee that we minimize the largest country deviation. Note that for each country $p$, exactly one of $\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ and $x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ is positive and exactly one is negative, unless both are zero. However, as soon as we reach $d_{t}^{*}\leq 1/2$ we have found a strongly close maximum cycle packing. Hence, in the remainder, we assume $d_{t}^{*}>1/2$ for each $t$ such that the series continues with $t+1$.
($\text{ILP}_{d_{1}}$) has one additional continuous variable ($d_{1}$) and $2|N|+1$ additional constraints. For every country $p\in N$, we have that $|\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}|\leq d_{1}^{*}$. However, there exists a smallest subset $N_{1}\subseteq N$ (which may not necessarily be unique) such that
$\begin{array}[]{ll}\displaystyle\left|\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}\right|=d_{1}^{*}&\forall p\in N_{1}\\ &\\ \displaystyle\left|\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}\right|\leq d_{2}^{*}<d_{1}^{*}&\forall p\in N\setminus N_{1}\end{array}$
@@
The number of vertices for a country covered by any cycle packing is an integer. Hence, the number of possible country deviations is at most $2|N|$ and depends only on $x$. The fractional part of a country deviation is either $\operatorname{frac}(x_{p})$ or $1-\operatorname{frac}(x_{p})$. Therefore, to find the minimal positive difference in between the deviations of any two countries $p$ and $q$, we have to compare the values $\operatorname{frac}(x_{p})$ and $1-\operatorname{frac}(x_{p})$, with $\operatorname{frac}(x_{q})$ or $1-\operatorname{frac}(x_{q})$ and take the minimum of those four possible differences. We let $\varepsilon$ be a small positive constant that is weakly smaller than the minimum possible positive difference between any two countries except in the unique case when $\operatorname{frac}(x_{p})=0.5$ for every $p\in N$. This is because in that case, $\operatorname{frac}(x_{p})=1-\operatorname{frac}(x_{p})=\operatorname{frac}(x_{q})=1-\operatorname{frac}(x_{q})=0.5$, meaning that all four possible differences are zero. However, should this happen, then the minimum possible positive difference of any two countries’ deviation can be trivially found as $1$, and then we can simply take any $\varepsilon$ with $0<\varepsilon\leq 1$.
@@
[…]
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $=$ $\displaystyle\sum_{j:ij\in A}e_{ij}$ $\quad\forall i\in V$ (1)
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $\leq$ $\quad 1$ $\quad\forall i\in V$ (2)
$\displaystyle\sum_{ij\in A}e_{ij}$ $=$ $\quad\mu^{*}$ (3)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad d_{1}^{*}-\varepsilon(1-z^{1}_{p})$ $\quad\forall p\in N$ (6)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad d_{1}^{*}-\varepsilon(1-z^{1}_{p})$ $\quad\forall p\in N$ (7)
@@
As discussed, for each country $p$, the left hand side of either (6) or (7) is negative (i.e., would be satisfied even with $z^{1}_{p}=0$). For those countries whose deviation cannot be lower than $d_{1}^{*}$, however, the (positive) left hand side of either (6) or (7) will require $z^{1}_{p}=1$. Thus, given an optimal solution $z^{1*}$ of ($\text{ILP}_{N_{1}}$), let $n_{1}^{*}:=\sum_{p\in N}z_{p}^{1*}$ be the minimal number of countries receiving the largest country deviations. It is guaranteed that the non-increasingly ordered country deviations at a strongly close maximum cycle packing starts with exactly $n_{1}^{*}$ many $d_{1}^{*}$ values, followed by some $d_{2}^{*}<d_{1}^{*}$. ($\text{ILP}_{N_{1}}$) has $|A|+|N|$ binary variables and $2|V|+2|N|+1$ constraints. Now, to find $d_{2}^{*}$, we solve the following ILP:
[…]
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $=$ $\displaystyle\sum_{j:ij\in A}e_{ij}$ $\quad\forall i\in V$ (1)
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $\leq$ $\quad 1$ $\quad\forall i\in V$ (2)
$\displaystyle\sum_{ij\in A}e_{ij}$ $=$ $\quad\mu^{*}$ (3)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad d_{1}^{*}$ $\quad\forall p\in N$ (8)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad d_{1}^{*}$ $\quad\forall p\in N$ (9)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad d_{2}+z^{1}_{p}d_{1}^{*}$ $\quad\forall p\in N$ (10)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad d_{2}+z^{1}_{p}d_{1}^{*}$ $\quad\forall p\in N$ (11)
$\displaystyle\sum_{p\in N}z^{1}_{p}$ $=$ $n_{1}^{*}$ (12)
@@
($\text{ILP}_{d_{2}}$) has $|A|+|N|$ binary variables and one continuous variable ($d_{2}$) with $2|V|+4|N|+2$ constraints, and guarantees that we find the minimal second-largest country deviation $d_{2}^{*}$ while exactly $n_{1}^{*}$ countries deviation is kept at $d_{1}^{*}$. Finding $n_{2}^{*}$ follows a similar approach, where $L$ is a large constant satisfying $L\geq 2(d_{1}^{*}-d_{2}^{*})$:
[…]
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $=$ $\displaystyle\sum_{j:ij\in A}e_{ij}$ $\quad\forall i\in V$ (1)
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $\leq$ $\quad 1$ $\quad\forall i\in V$ (2)
$\displaystyle\sum_{ij\in A}e_{ij}$ $=$ $\quad\mu^{*}$ (3)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L$ $\quad\forall p\in N$ (13)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L$ $\quad\forall p\in N$ (14)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad d_{1}^{*}$ $\quad\forall p\in N$ (8)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad d_{1}^{*}$ $\quad\forall p\in N$ (9)
$\displaystyle\sum_{p\in N}z^{1}_{p}$ $=$ $\quad n_{1}^{*}$ (12)
$z_{p}^{1}+z_{p}^{2}$ $\leq$ $1$ $\quad\forall p\in N$ (15)
@@
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $=$ $\displaystyle\sum_{j:ij\in A}e_{ij}$ $\quad\forall i\in V$ (1)
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $\leq$ $\quad 1$ $\quad\forall i\in V$ (2)
$\displaystyle\sum_{ij\in A}e_{ij}$ $=$ $\quad\mu^{*}$ (3)
$\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}$ $\leq$ $\quad 1$ $\quad\forall p\in N$ (16)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad d_{t}+\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}$ $\quad\forall p\in N$ (17)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad d_{t}+\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}$ $\quad\forall p\in N$ (18)
$\displaystyle\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}-x_{p}$ $\leq$ $\quad\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}+{\left(1-\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}\right)}L$ $\quad\forall p\in N$ (19)
$\displaystyle x_{p}-\sum_{j\in V_{p}}{\sum_{i:ij\in A}}e_{ij}$ $\leq$ $\quad\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}+{\left(1-\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}\right)}L$ $\quad\forall p\in N$ (20)
$\displaystyle\sum_{p\in N}z^{i}_{p}$ $=$ $\quad n_{i}^{*}$ $\quad\forall{i\in\{1,\dots,t-1\}}$ (21)
In the following formulation $L^{\prime}$ is a large constant satisfying ${L^{\prime}}\geq d_{t}^{*}-\varepsilon$.
[…]
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $=$ $\displaystyle\sum_{j:ij\in A}e_{ij}$ $\quad\forall i\in V$ (1)
$\displaystyle\sum_{j:ji\in A}e_{ji}$ $\leq$ $\quad 1$ $\quad\forall i\in V$ (2)
$\displaystyle\sum_{ij\in A}e_{ij}$ $=$ $\quad\mu^{*}$ (3)
$\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}$ $\leq$ $\quad 1$ $\quad\forall p\in N$ (16)
$\displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}$ $\leq$ $\quad d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}$ $\quad\forall p\in N$ (22)
$x_{p}-\displaystyle\sum_{j\in V_{p}}e_{ij}$ $\leq$ $\quad d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}$ $\quad\forall p\in N$ (23)
$\displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}$ $\leq$ $\quad\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+{\left(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i}\right)}{L^{\prime}}$ $\quad\forall p\in N$ (24)
$\displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}$ $\leq$ $\quad\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+{\left(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i}\right)}{L^{\prime}}$ $\quad\forall p\in N$ (25)
$\displaystyle\sum_{p\in N}z^{i}_{p}$ $=$ $\quad n_{i}^{*}$ $\quad\forall{i\in\{1,\dots,t-1\}}$ (21)
From the above, we conclude that the following theorem holds, (in which the bounds on the number of ILPs, variables and constraints are not necessarily tight).
@@
For a partitioned permutation game $(N,v)$ defined on a graph $G=(V,A)$, it is possible to find an optimal solution that is strongly close to a given target allocation $x$ by solving a series of at most $2|N|$ ILPs, each having $O(|A|+|N|^{2})$ binary variables and $O(|V|)$ constraints.